Circular No. 8113

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PROBABLE NOVA IN SAGITTARIUS S. Nakano, Sumoto, Japan, reports the discovery by Hideo Nishimura

(Kakegawa, Shizuoka-ken) of a possible nova (mag 9.8) on a T-Max 400 film taken on Apr. 8.792 UT with a 200-mm f/4 camera lens. Nishimura reported the object at α ≈ 18^{h}07^{m}19^{s}, δ ≈ −27^{o}24′20′′ (equinox 2000.0), and he provided the following magnitudes from other films: Mar. 25.81, [11; 27.81, 10.9; 30.80, [11–12; Apr. 5.80, 10.4. R. Kushida, Yatsugatake South Base Observatory, reports that an unfiltered CCD frame taken on Apr. 9.704 shows the object at mag 7.6 and yields the following position end figures (measured by Y. Kushida): 20^{s}.38, 31′′.6. She adds that nothing appears at this position brighter than mag 18–19 on the Digitized Sky Survey (two stars of mag 19–20 appear near the position).

SUPERNOVA 2003dg IN UGC 6934 Further to IAUC 8101, H. Pugh and W. Li report the LOTOSS dis-

covery of an apparent supernova on unfiltered KAIT images on Apr. 8.4 (mag ∼ 19.3) and 9.4 UT (mag ∼ 19.1). The new object is located at α = 11^{h}57^{m}31^{s}.97, δ = −1^{o}15′13′′.6 (equinox 2000.0), which is 2′′.7 east and 3′′.1 south of the nucleus of UGC 6934. A KAIT image taken on Mar. 24.4 showed nothing at this position (limiting mag ∼ 20.0).

XTE J1807–294 S. Campana and M. Ravasio, Osservatorio Astronomico di Brera; and

G. L. Israel and V. Mangano, Osservatorio Astronomico di Roma, report: “XMM-Newton observed the fourth millisecond pulsar XTE J1807–294 (IAUC 8080) during Mar. 22.581–22.778 UT. The source position is consistent with that determined with Chandra (IAUC 8095). The 0.5–10-keV x-ray countrate measured with EPIC-pn (in timing mode) was ≈ 34.2 ± 0.1 count/s. No x-ray bursts were detected. Pulsations at 5.2459996(4) ms were detected with a pulsed fraction of 1.9 ± 0.3 percent (90-percent confidence level) in the band 0.5–10 keV. The orbital modulation instead was barely visible (probably due to the short exposure). No prominent emission or absorption lines are present in the RGS spectra. Analysis of the EPIC-MOS1 (full-frame mode) and EPIC-pn (timing mode) provided an absorbed (0.5–10 keV) source flux of 2 × 10^{−10} erg cm^{−2} s^{−1}. The source is still bright, and optical and infrared observations are encouraged in order to identify the optical counterpart.”

2003 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green